# MIMDRAM: An End-to-End Processing-Using-DRAM System for High-Throughput, Energy-Efficient and Programmer-Transparent Multiple-Instruction Multiple-Data Processing

### 29 Feb 2024 v1

@@ 1 Introduction
Data movement between computation units (e.g., CPUs, GPUs) and main memory (e.g., DRAM) is a major performance and energy bottleneck in current computing systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], and is expected to worsen due to the increasing data intensiveness of modern applications [21, 22]. To mitigate the overheads caused by data movement, several works propose processing-in-memory (PIM) architectures [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 17, 10, 38, 39, 40, 41, 42, 18, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117]. There are two main approaches to PIM [8, 9]: (i) processing-near-memory (PNM) [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 17, 10, 38, 39, 40, 41, 42, 18, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 104], where computation logic is added near the memory arrays (e.g., in a DRAM chip or at the logic layer of a 3D-stacked memory [118, 119, 120]); and (ii) processing-using-memory (PUM) [59, 60, 61, 62, 63, 64, 65, 66, 67, 121, 68, 69, 70, 71, 72, 52, 73, 74, 75, 76, 77, 78, 79, 122, 123, 80, 81, 82, 83, 124, 84, 85, 86, 87, 125, 88, 126, 127, 128, 89, 90, 129, 91, 92, 93, 94, 95, 130, 131, 132, 133, 134, 135, 136, 137, 138, 96, 139, 140, 97, 141, 142, 143, 98, 99, 144, 100, 101, 145, 102, 103, 105, 146], where computation is performed by exploiting the analog operational properties of the memory circuitry. There are two main advantages of PUM over PNM. First, PUM fundamentally eliminates data movement by performing computation in situ, while data movement still occurs between computation units and memory arrays in PNM. Second, PUM architectures exploit the large internal bandwidth and parallelism available inside the memory arrays, while PNM solutions are fundamentally bottlenecked by the memory’s internal data buses.
@@
PUM architectures can be implemented using different memory technologies, including SRAM [71, 70, 96, 97, 98, 99], DRAM [125, 52, 85, 66, 78, 79, 69, 63, 88, 83, 86, 74, 64, 76, 73, 65, 61, 75, 62, 67], emerging non-volatile [77, 60, 68, 82, 81, 90, 59] or NAND flash [80, 91, 92, 93, 94, 95]. Processing-using-DRAM (PUD) [52, 85, 66, 78, 79, 69, 63, 88, 83, 86, 74, 64, 76, 73, 65, 61, 75, 62, 67], in particular, enables the execution of different bulk operations in DRAM (i.e., PUD operations), such as (i) data copy and initialization [74, 64, 75, 100], (ii) bitwise Boolean operations [61, 69, 67, 52, 63], (iii) arithmetic operations [66, 69, 63, 125, 101, 88], and (iv) lookup table based operations [78, 79, 145, 102, 103].

### 03 Mar 2024 v2

@@ 1 Introduction
Data movement between computation units (e.g., CPUs, GPUs) and main memory (e.g., DRAM) is a major performance and energy bottleneck in current computing systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], and is expected to worsen due to the increasing data intensiveness of modern applications [21, 22]. To mitigate the overheads caused by data movement, several works propose processing-in-memory (PIM) architectures [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 17, 49, 10, 50, 51, 52, 53, 54, 55, 56, 18, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189, 190]. There are two main approaches to PIM [8, 9]: (i) processing-near-memory (PNM) [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 17, 49, 10, 50, 51, 52, 53, 54, 55, 56, 18, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 176], where computation logic is added near the memory arrays (e.g., in a DRAM chip or at the logic layer of a 3D-stacked memory [191, 192, 193]); and (ii) processing-using-memory (PUM) [106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 73, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 177, 183], where computation is performed by exploiting the analog operational properties of the memory circuitry. There are two main advantages of PUM over PNM. First, PUM fundamentally eliminates data movement by performing computation in situ, while data movement still occurs between computation units and memory arrays in PNM. Second, PUM architectures exploit the large internal bandwidth and parallelism available inside the memory arrays, while PNM solutions are fundamentally bottlenecked by the memory’s internal data buses.
@@
PUM architectures can be implemented using different memory technologies, including SRAM [119, 118, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169], DRAM [139, 73, 136, 113, 126, 127, 117, 110, 140, 133, 137, 122, 111, 124, 121, 112, 108, 123, 109, 114], emerging non-volatile [128, 125, 107, 116, 115, 132, 131, 145, 106, 170] or NAND flash [130, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158]. Processing-using-DRAM (PUD) [139, 73, 136, 113, 126, 127, 117, 110, 140, 133, 137, 122, 111, 124, 121, 112, 108, 123, 109, 114], in particular, enables the execution of different bulk operations in DRAM (i.e., PUD operations), such as (i) data copy and initialization [122, 111, 123, 171], (ii) bitwise Boolean operations [108, 117, 114, 73, 110], (iii) arithmetic operations [113, 117, 110, 139, 172, 140], and (iv) lookup table based operations [126, 127, 173, 174, 175].